OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 3 and contains the point (4, 9)

OpenStudy (mathstudent55):

The slope-intercept form of the equation of a line is: y = mx + b where m = slope and b = the y-intercept. You have the slope = m = 3, so let's replace m with 3: y = 3x + b To find b, insert the known point (x, y) = (4, 9) into the x and y variables of the equation and solve for b. Then replace b in y = 3x + b with the value of b you find.

OpenStudy (anonymous):

Slope intercept form is:y=mx+b m=3 given since the slope is =rise/run=change in y divided by change in x ,m=3 tells us that for a change of 1 in x we have a change of 3 in y one point is given (4,9) substituting back into the equation we have 9=3(4)+b =>9-12=b =>-3=b so,the intercept appears to be -3 the resulting equation is y=3x-3